# Cross-Problem Solving for Network Optimization: Is Problem-Aware Learning the Key?

### v1

@@ Abstract
As intelligent network services continue to diversify, ensuring efficient and adaptive resource allocation in edge networks has become increasingly critical. Yet the wide functional variations across services often give rise to new and unforeseen optimization problems, rendering traditional manual modeling and solver design both time-consuming and inflexible. This limitation reveals a key gap between current methods and human solving — the inability to recognize and understand problem characteristics. It raises the question of whether problem-aware learning can bridge this gap and support effective cross-problem generalization. To answer this question, we propose a problem-aware diffusion (PAD) model, which leverages a problem-aware learning framework to enable cross-problem generalization. By explicitly encoding the mathematical formulations of optimization problems into token-level embeddings, PAD empowers the model to understand and adapt to problem structures. Extensive experiments across six diverse network optimization problems show that PAD generalizes well to unseen problems while significantly improving solution quality and feasibility. Meanwhile, an auxiliary constraint-aware module is designed to enforce solution validity further. The experiments reveal that problem-aware learning is promising for building general-purpose solvers for intelligent network operation and resource management. Our code is open source at https://github.com/qiyu3816/PAD.
Index Terms:
Network optimization, resource allocation, problem-aware learning, generative diffusion model.
@@ I Introduction
The vision of future wireless communication systems has been widely built in both academia and industry [1, 2, 3, 4, 5], with fundamental AI models capable of general problem-solving being pursued as a revolutionary technology. These intelligent fundamental AI models are expected to enable autonomous and general network operations across various tasks, such as network resource management [6] and data analysis [7]. Notably, the continuous emergence of novel intelligent services in wireless networks is driving an unprecedented demand for efficient allocation of network resources like power and spectrum. In future wireless networks, different types of tasks—such as power allocation and spectrum allocation—will coexist within edge networks, while edge networks themselves may differ significantly in functionality and environmental settings (e.g., computation offloading networks and edge sensing networks), which inherently raises diverse network optimization problems with varying objectives, constraints, and mathematical formulations [5, 8, 9, 10, 11]. However, most existing network optimization approaches treat each problem in isolation, failing to address them in a unified and seamless manner. Consequently, cross-problem network optimization is a key capability for the intelligent fundamental models of next-generation wireless communication systems.
@@
In this context, a key question raised: Can problem-aware learning enable models to develop a meaningful understanding of problem objectives and achieve cross-problem generalization? Inspired by recent advances in using large language models (LLMs) for mathematical reasoning [22], math question similarity calculation [23], and optimization problem modeling [24], we argue that the mathematical formulations of network optimization problems also contain instructive information that can help models better understand problem characteristics. Incorporating these representations into the model’s inference process can enhance the capacity for cross-problem solving. To address the aforementioned challenges, we propose a problem-aware diffusion (PAD) model designed for cross-problem solution generation in network optimization within a single fundamental model. PAD explores a novel problem-aware approach that integrates the rich mathematical representations of optimization problems into model learning. To achieve this, PAD leverages feature embeddings of network optimization problem formulations from existing mathematical LLMs during both training and solution generation.
@@
  • Can a constraint-aware module improve the feasibility of generated solutions? PAD introduces a constraint-aware module, which significantly improves constraint satisfaction by incorporating a constraint loss during training. This provides an efficient mechanism for ensuring that the model’s outputs comply with the hard constraints of the original optimization problems.
  • How can raw problem-specific feature embeddings be effectively utilized in problem-aware learning? To effectively extract key information from the token sequence of the raw problem embedding, we design a ranked pooling mechanism based on intra-sequence token similarity scores. This approach not only reduces sequence length and removes irrelevant information, but also remains highly sensitive to subtle changes in problem semantics. Empirically, our method outperforms other standard pooling techniques.
@@ II-A Solution Generation for Network Optimization
Solution generation methods based on LLMs [29, 9, 30] and generative diffusion models (GDMs) [6, 21, 31, 32, 33] have achieved remarkable success in network optimization tasks. These approaches leverage the strong logical reasoning capabilities of LLMs and the powerful solution distribution modeling of diffusion models to produce superior solutions for specific problems. As data-driven approaches, generative models differ fundamentally from discriminative models. While discriminative models learn conditional probability distributions, generative models learn joint distributions, which makes their inference process inherently a form of sampling rather than classification or regression. This sampling-based inference allows for parallel generation and offers a more global understanding of the task compared to the fixed-output nature of common regression models. Generative AI (GenAI)-based frameworks for network management and optimization are rapidly gaining momentum in the communication network community. With their strong scaling laws, GenAI is increasingly seen as a promising candidate for serving as a foundational model in future communication systems.
Figure 1: Comparison of PAD with other isolated methods for cross-problem network optimization. PAD trains once and adapts to new problems via explicitly input $\mathbb{P}_{i}$, while traditional methods require re-training, even re-designing, for each individual problem.
@@ II-C Problem-Aware Learning
Problem-aware learning refers to a class of learning methods being explicitly informed or guided by the structure, objectives, or constraints of the underlying problem being solved, instead of learning patterns purely from data in a black-box way [36]. For example, [37] employs LLMs to iteratively generate solutions for simple optimization problems based on natural language descriptions, and [30] leverages domain-specific LLMs to produce resource allocation solutions for network optimization problems described in natural language. In addition, [38] advocates integrating formal methods (such as mathematical algorithms or rules) with machine learning, similar to problem-aware learning. However, [38] is limited to combining simple constrained problems into losses in the form of Lagrange functions, and is challenging to apply to network optimization problems where non-differentiable functions are standard.
@@
At the same time, the potential of GDMs for cross-problem generalization has been validated in the AI community [39, 40]. Nevertheless, these works primarily focus on control problems, which differ significantly from network optimization in both formulation and application. Therefore, we believe integrating GDMs with problem-aware learning is a promising direction toward achieving cross-problem solving for network optimization.
@@ III-A Cross-Problem Network Optimization
where $M$-dimensional vector $\mathbf{x}\in\mathbb{R}^{M}$ denotes the input parameters such as network states and task-related information, while $N$-dimensional vector $\mathbf{y}\in\mathbb{R}^{N}$ represents the optimization variables such as power allocation and spectrum assignment. The functions $f:\mathbb{R}^{M+N}\rightarrow\mathbb{R}$ and $g:\mathbb{R}^{M+N}\rightarrow\mathbb{R}$ correspond to the objective function and the constraint function, respectively.
Cross-problem network optimization refers to the ability of a method to generalize across structurally and semantically different optimization problems, while preserving solution quality and not requiring task-specific redesign. The problems in this set may differ in objective function structure, constraint form, and semantic interpretation of inputs and outputs. As illustrated in Fig. 1, different network optimization problems such as $\mathbb{P}_{1}$ and $\mathbb{P}_{2}$ may have distinct mathematical semantics in their objective functions $f_{1}$ and $f_{2}$, as well as in their constraint functions $g_{1}$ and $g_{2}$. They may also differ in their input parameters $\mathbf{x}_{1},\mathbf{x}_{2}$ and output variables $\mathbf{y}_{1},\mathbf{y}_{2}$.
Traditional isolated optimizers rely on intensive human effort while limiting their applicability to narrowly scoped scenarios. Cross-problem network optimization offers a promising path toward efficient, automated, and scalable network operations in increasingly dynamic environments. However, as described in Section II, these models do not explicitly incorporate the mathematical structure of the problems during training, resulting in limited adaptability to new tasks.
@@ III-B Selected Optimization Problems
In this section, we provide an overview of the set of network optimization problems considered in our implementation. The problem selection is based on three key principles: 1) the problems must involve the allocation of major network resources; 2) the problems should include diverse types of input features and optimization variables with different physical meanings; 3) the problems should cover a wide range of objectives and constraints to reflect the diversity of optimization problems encountered in real-world scenarios. Specifically, we consider six problems:
@@
  3. 3.
    Power allocation aiming to minimize the maximum transmission delay;
  4. 4.
  5. 5.
@@
    Spectrum resource block allocation aiming to maximize total transmission rate, subject to fairness constraints.
@@
These problems include two fundamental resource types: power [30, 41, 42] and spectrum [31, 43, 44], which are among the most critical optimization variables in wireless networks. The selected problems involve both continuous (real-valued) and discrete (integer-valued) variables, reflecting the real-world diversity of decision spaces. Standard parameters such as channel gains, data volumes, and noise power are incorporated to enhance the realism further. The objectives and constraints of these problems are inspired by key application scenarios, including multi-access edge computing (MEC) networks [45], integrated sensing and communication (ISAC) networks [3], and federated learning (FL) systems [46]. The problems cover various optimization types, from classical convex optimization to more challenging non-convex optimization and integer programming. By evaluating the proposed PAD across this selected and diverse problem set, we aim to demonstrate its effectiveness in handling different optimization structures and generalizing across diverse tasks.
@@ IV-A Basics
$p(\mathbf{y}_{i}|\mathbf{e}_{i},\mathbf{x}_{i})=\frac{p(\mathbf{y}_{i},\mathbf{e}_{i},\mathbf{x}_{i})}{p(\mathbf{e}_{i},\mathbf{x}_{i})},$ (3)
In the context of network optimization, different problems may share the same network state parameters and resource allocation variables. As a result, the optimization variable $\mathbf{y}$ and the input parameter $\mathbf{x}$ are both conditionally independent of the problem embedding $\mathbf{e}$. Therefore, we have $p(\mathbf{e}_{i},\mathbf{x}_{i})=p(\mathbf{e}_{i})p(\mathbf{x}_{i})$, where $p(\mathbf{e}_{i})<1$, leading to
$p(\mathbf{y}_{i}|\mathbf{e}_{i},\mathbf{x}_{i})>p(\mathbf{y}_{i},\mathbf{e}_{i}|\mathbf{x}_{i}).$ (4)
Intuitively, Eq. (4) implies that the learning objective in Eq. (3) is more deterministic and effective than that in Eq. (2).
Specifically, due to the independence between $\mathbf{x}$ and $\mathbf{e}$, attempting to infer the problem embedding $\mathbf{e}_{i}$ directly from $\mathbf{x}_{i}$ in Eq. (2) is theoretically flawed and likely to introduce noise. Moreover, assuming the output space of $\mathbf{y}$ is of size $Y$, and that of $\mathbf{e}$ is $E$ (even though both are infinite in practice), the output space of Eq. (2) becomes $Y\times E$, which is significantly larger than the output space $Y$ of Eq. (3). This leads to a sparsity of training samples per output configuration in Eq. (2), making learning less efficient and more prone to underfitting.
@@
IV-B Ranked Pool for Problem Feature Embedding
@@
    Structural Sparsity. Mathematical formulations of optimization problems are inherently structured, featuring segmented semantic components such as objective functions, constraints, and optimization variables. However, the corresponding tokens are often distributed sparsely and non-contiguously across the sequence, making retrieving meaningful structure more difficult.
@@
    Symbol Sensitivity. Small changes in mathematical symbols can lead to entirely different problem types. Moreover, global operators such as $\mathrm{min}$, $\sum$, or nested formulations require the structural integrity of the expression to be preserved during token selection.
@@
    Long-Tail Distribution. Standard low-level tokens (e.g., basic operators or variable names) appear frequently but carry limited semantic value. In contrast, high-information global symbols appear less often but are critical for understanding problem structure. A robust token selection mechanism must retain and emphasize these rare but informative tokens.
Conventional pooling strategies struggle to handle these properties. For instance, MaxPool [47], which selects the most active tokens at each dimension, is highly susceptible to noise and often overemphasizes frequent but low-informative symbols (e.g., repeating input variable names), while potentially breaking mathematical structure by discarding crucial operator relationships. AvgPool [47], on the other hand, dilutes the semantic strength of key tokens and is insensitive to positional structure, failing to preserve the underlying formulation hierarchy. Although these methods have been widely adopted in other tasks, they demonstrate poor performance when applied to token selection in problem embeddings due to the unique structural and symbolic demands of mathematical optimization representations.
We propose a token selection method based on intra-sequence token similarity scores derived from the problem embedding, referred to as the ranked pool. Specifically, given a raw problem embedding sequence of length $S$, we define a similarity score matrix $\mathbf{M}_{\mathrm{SI}}$, where each entry quantifies the similarity between a pair of token vectors. We compute pairwise similarities using PyTorch’s built-in cosine_similarity() function [47], where values approaching 1 indicate high similarity and those near 0 indicate low similarity. Formally, the similarity matrix is defined as
@@
By selecting the most diverse tokens, the ranked pool effectively captures unique and informative symbolic elements, especially for those sensitive to small changes in mathematical expressions. Additionally, keeping the selected tokens in their original order helps preserve important structural relationships within the expression. In practice, ranked pool can be performed only once when the problem is first encountered, and the resulting pooled sequence can be cached for reuse, thereby avoiding repeated calls to the LLM for encoding and pooling. The experimental details are shown in Section V-B.
Input: $N$ training problems and their corresponding datasets $\boldsymbol{\mathcal{D}}=\{\mathcal{D}_{1},\ldots,\mathcal{D}_{i},\ldots,\mathcal{D}_{N}\}$, the encoder model $\boldsymbol{\theta}_{\mathrm{E}}$, the decoder model $\boldsymbol{\theta}_{\mathrm{D}}$;
@@
5                   $\mathbf{x},\mathbf{y}$leftarrow batch data from $\mathcal{D}_{i}$;
6                   $\mathbf{v}_{\mathrm{joint}}$leftarrow cat($\mathbf{x},\mathbf{y}$);
7                   $\tilde{\mathbf{v}}_{\mathrm{joint}}$leftarrow $f_{\boldsymbol{\theta}_{\mathrm{D}}}(f_{\boldsymbol{\theta}_{\mathrm{E}}}(\mathbf{x}))$;
8                   $\mathbf{v}_{\mathrm{val}},\mathbf{v}_{\mathrm{type}}$leftarrow split $\mathbf{v}_{\mathrm{joint}}$;
9                   $\tilde{\mathbf{v}}_{\mathrm{val}},\tilde{\mathbf{v}}_{\mathrm{type}}$leftarrow split $\tilde{\mathbf{v}}_{\mathrm{joint}}$;
10                   $\mathcal{L}_{\mathrm{real}}$leftarrow $\|\mathbf{v}_{\mathrm{val}}-\tilde{\mathbf{v}}_{\mathrm{val}}\|^{2}$;
11                   $\mathcal{L}_{\mathrm{class}}$leftarrow $\mathrm{CrossEntropyLoss}(\mathbf{v}_{\mathrm{type}},\tilde{\mathbf{v}}_{\mathrm{type}})$;
12                   $\mathcal{L}$leftarrow $\mathcal{L}_{\mathrm{real}}+\mathcal{L}_{\mathrm{class}}$;

### v2

@@ Abstract
As intelligent network services continue to diversify, ensuring efficient and adaptive resource allocation in edge networks has become increasingly critical. Yet the wide functional variations across services often give rise to new and unforeseen optimization problems, rendering traditional manual modeling and solver design both time-consuming and inflexible. This limitation reveals a key gap between current methods and human solving — the inability to recognize and understand problem characteristics. It raises the question of whether problem-aware learning can bridge this gap and support effective cross-problem generalization. To answer this question, we propose a problem-aware diffusion (PAD) model, which leverages a problem-aware learning framework to enable cross-problem generalization. By explicitly encoding the mathematical formulations of optimization problems into token-level embeddings, PAD empowers the model to understand and adapt to problem structures. Extensive experiments across ten representative network optimization problems show that PAD generalizes well to unseen problems while avoiding the inefficiency of building new solvers from scratch, yet still delivering competitive solution quality. Meanwhile, an auxiliary constraint-aware module is designed to enforce solution validity further. The experiments indicate that problem-aware learning opens a promising direction toward general-purpose solvers for intelligent network operation and resource management. Our code is open source at https://github.com/qiyu3816/PAD.
@@ I Introduction
The vision of future wireless communication systems has been widely built in both academia and industry [1, 2, 3, 4, 5], with fundamental AI models capable of general problem-solving being pursued as a revolutionary technology. These intelligent fundamental AI models are expected to enable autonomous and general network operations across various tasks, such as network resource management [6] and data analysis [7]. Notably, the continuous emergence of novel intelligent services in wireless networks is driving an unprecedented demand for efficient allocation of network resources like power and spectrum. In future wireless networks, different types of tasks—such as power allocation and spectrum allocation—will coexist within edge networks, while edge networks themselves may differ significantly in functionality and environmental settings (e.g., computation offloading networks and edge sensing networks), which inherently raises diverse network optimization problems with varying objectives, constraints, and mathematical formulations [5, 8, 9, 10, 11, 12]. However, most existing network optimization approaches treat each problem in isolation, failing to address them in a unified and seamless manner. Consequently, cross-problem network optimization is a key capability for the intelligent fundamental models of next-generation wireless communication systems.
@@
In this context, a key question raised: Can problem-aware learning enable models to develop a meaningful understanding of problem objectives and achieve cross-problem generalization? Inspired by recent advances in using large language models (LLMs) for mathematical reasoning [23], math question similarity calculation [24], and optimization problem modeling [25], we argue that the mathematical formulations of network optimization problems also contain instructive information that can help models better understand problem characteristics. Incorporating these representations into the model’s inference process can enhance the capacity for cross-problem solving. To address the aforementioned challenges, we propose PAD, a problem-aware diffusion model designed for cross-problem solution generation in network optimization within a single fundamental model. PAD explores a novel problem-aware approach that integrates the rich mathematical representations of optimization problems into model learning. To achieve this, PAD leverages feature embeddings of network optimization problem formulations from existing mathematical LLMs during both training and solution generation.
@@
  • How can raw problem-specific feature embeddings be effectively utilized in problem-aware learning? To effectively extract key information from the token sequence of the raw problem embedding, we design a rank pool mechanism based on intra-sequence token similarity scores. This approach not only extracts essential information from the raw embedding and remains sensitive to structural variations, but also provides greater potential to capture the most distinctive features that differentiate one problem from others. Empirically, our method outperforms other standard pooling techniques.
  • Can a constraint-aware module improve the feasibility of generated solutions? PAD introduces a constraint-aware module, which markedly improves constraint satisfaction by incorporating a constraint loss during training. This provides an efficient mechanism for ensuring that the model’s outputs comply with the hard constraints of the original optimization problems.
@@ II-A Solution Generation for Network Optimization
Solution generation methods based on LLMs [30, 9, 31] and generative diffusion models (GDMs) [6, 22, 32, 33, 34] have achieved remarkable success in network optimization tasks. These approaches leverage the strong logical reasoning capabilities of LLMs and the powerful solution distribution modeling of diffusion models to produce superior solutions for specific problems. As data-driven approaches, generative models differ fundamentally from discriminative models. While discriminative models learn conditional probability distributions, generative models learn joint distributions, which makes their inference process inherently a form of sampling rather than classification or regression. This sampling-based inference enables parallel generation and provides a more comprehensive understanding of the task compared to the fixed-output nature of conventional regression models. Generative AI (GenAI)-based frameworks for network management and optimization are rapidly gaining momentum in the communication network community. With their strong scaling laws, GenAI is increasingly seen as a promising candidate for serving as a foundational model in future communication systems.
Figure 1: Comparison of PAD with other isolated methods for cross-problem network optimization. PAD trains once and adapts to new problems by explicitly inputting $\mathbb{P}_{i}$ from formulation embedding, whereas traditional methods require re-training, even redesigning, for each individual problem.
@@ II-C Problem-Aware Learning
Problem-aware learning refers to a class of learning methods being explicitly informed or guided by the structure, objectives, or constraints of the underlying problem being solved, instead of learning patterns purely from data in a black-box way [37]. For example, [38] employs LLMs to iteratively generate solutions for simple optimization problems based on natural language descriptions, and [31] leverages domain-specific LLMs to produce resource allocation solutions for network optimization problems described in natural language. In addition, [39] advocates integrating formal methods (such as mathematical algorithms or rules) with machine learning, similar to problem-aware learning. However, [39] is limited to combining simple constrained problems into losses in the form of Lagrange functions, and is challenging to apply to network optimization problems where non-differentiable functions are frequently encountered.
@@
At the same time, the potential of GDMs for cross-problem generalization has been validated in the AI community [40, 41]. Nevertheless, these works primarily focus on automatic control problems, which differ significantly from network optimization in both formulation and application. Therefore, we believe that integrating GDMs with problem-aware learning is a promising direction for achieving cross-problem solving in network optimization.
@@ III-A Cross-Problem Network Optimization
where $M$-dimensional vector $\mathbf{x}\in\mathbb{R}^{M}$ denotes the input parameters, such as network states and task-related information, while $N$-dimensional vector $\mathbf{y}\in\mathbb{R}^{N}$ represents the optimization variables, such as power allocation and spectrum assignment. The functions $f:\mathbb{R}^{M+N}\rightarrow\mathbb{R}$ and $g:\mathbb{R}^{M+N}\rightarrow\mathbb{R}$ correspond to the objective function and the constraint function, respectively. In network optimization, $f$ typically corresponds to global objectives that need to be maximized or minimized, such as transmission rate, throughput, or resource cost. $g$, on the other hand, is usually used to enforce physical resource constraints and prevent logical conflicts.
Cross-problem network optimization refers to the ability of a method to generalize across structurally and semantically different optimization problems, while preserving solution quality and not requiring task-specific redesign. The problems in this set may differ in terms of objective function structure, constraint form, and the semantic interpretation of inputs and outputs. As illustrated in Fig. 1, different network optimization problems such as $\mathbb{P}_{1}$ and $\mathbb{P}_{2}$ may have distinct mathematical semantics in their objective functions $f_{1}$ and $f_{2}$, as well as in their constraint functions $g_{1}$ and $g_{2}$. They may also differ in their input parameters $\mathbf{x}_{1},\mathbf{x}_{2}$ and output variables $\mathbf{y}_{1},\mathbf{y}_{2}$.
Traditional isolated optimizers rely heavily on intensive human effort, limiting their applicability to narrowly scoped scenarios. Cross-problem network optimization presents a promising approach to efficient, automated, and scalable network operations in increasingly dynamic environments. However, as described in Section II, these models do not explicitly incorporate the mathematical structure of the problems during training, resulting in limited adaptability to new tasks.
@@ III-B Selected Optimization Problems
In this section, we provide an overview of the set of network optimization problems considered in our implementation. The problem selection is based on three key principles: 1) the problems must involve the allocation of major network resources; 2) the problems should include diverse types of input features and optimization variables with different physical meanings; 3) the problems should cover a wide range of objectives and constraints to reflect the diversity of optimization problems encountered in real-world scenarios. Specifically, we consider the following ten problems:
@@
  3. 3.
    Spectrum resource block allocation aiming to maximize total transmission rate, subject to fairness constraints;
  4. 4.
  5. 5.
@@
    Cache refresh rate optimization aiming to minimize expected hit latency;
  7. 7.
    Joint spectrum and power allocation aiming to maximize total transmission rate, subject to minimum rate constraints per link;
  8. 8.
    Cache decision optimization aiming to minimize expected hit latency;
  9. 9.
    Joint offloading decision and computational resource optimization aiming to minimize overall task offloading cost;
  10. 10.
    Power allocation aiming to minimize the maximum transmission delay.
@@
These problems include several common resource types: power [31, 42, 43], spectrum [32, 44, 45, 46], caching resources [47, 48, 49], and computational resources [50, 51], which together represent the core decision variables in modern network optimization. The selected problems involve both continuous (real-valued) and discrete (integer-valued) variables, reflecting the real-world diversity of decision spaces. Standard parameters, such as channel gains, data volumes, and noise power, are incorporated to further enhance the realism. The objectives and constraints of these problems are inspired by key application scenarios, including multi-access edge computing (MEC) networks [52, 50], integrated sensing and communication (ISAC) networks [3, 53], federated learning (FL) systems [54], and contaimed at minimizing the maximum transmission delay, is formulated as a bi-level optimization problem, which is inherently more challenging and is a hematical optimization type encountered in network optimization. Specifically, the first nine problems are grouped into three clusters, each containing three problems and corresponding to a representative optimization paradigm: convex optimization, non-convex optimization, and mixed-integer or combinatorial optimization. The final problem, power allocation aiming to minimize the maximum transmission delay, is formulated as a bi-level optimization problem, which is inherently more challenging and therefore presented separately. In the following model design, only one representative problem from each cluster is used for training, while the remaining problems are reserved for testing and evaluation of generalization. By evaluating the proposed PAD across this selected and diverse problem set, we aim to demonstrate its effectiveness in handling different optimization structures and generalizing across diverse tasks.
These problems include several common resource types: power [31, 42, 43], spectrum [32, 44, 45, 46], caching resources [47, 48, 49], and computational resources [50, 51], which together represent the core decision variables in modern network optimization. The selected problems involve both continuous (real-valued) and discrete (integer-valued) variables, reflecting the real-world diversity of decision spaces. Standard parameters, such as channel gains, data volumes, and noise power, are incorporated to further enhance the realism. The objectives and constraints of these problems are inspired by key application scenarios, including multi-access edge computing (MEC) networks [52, 50], integrated sensing and communication (ISAC) networks [3, 53], federated learning (FL) systems [54], and content caching and delivery (CDN) networks [48, 49].
The above problems cover the common mathematical optimization types encountered in network optimization. Specifically, the first nine problems are grouped into three clusters, each containing three problems and corresponding to a representative optimization paradigm: convex optimization, non-convex optimization, and mixed-integer or combinatorial optimization. The final problem, power allocation to minimize the maximum transmission delay, is formulated as a bi-level optimization problem and is inherently more challenging; therefore, it is presented separately. In the following model design, only one representative problem from each cluster is used for training, while the remaining problems are reserved for testing and evaluation of generalization. By evaluating the proposed PAD across this selected and diverse problem set, we aim to demonstrate its effectiveness in handling different optimization structures and generalizing across diverse tasks.
@@ IV-A Basics
$p(\mathbf{y}_{i}|\mathbf{e}_{i},\mathbf{x}_{i})=\frac{p(\mathbf{y}_{i},\mathbf{e}_{i},\mathbf{x}_{i})}{p(\mathbf{e}_{i},\mathbf{x}_{i})}.$ (3)
In the context of network optimization, different problems may share the same network state parameters and resource allocation variables, yet certain specialized parameters and variables may correspond exclusively to specific problems. As a result, the input parameter $\mathbf{x}$ and the problem embedding $\mathbf{e}$ are not conditionally independent, and their correspondence is inherently non-deterministic. Therefore, we have
$p(\mathbf{e}_{i},\mathbf{x}_{i})=p(\mathbf{e}_{i}|\mathbf{x}_{i})p(\mathbf{x}_{i})\leq p(\mathbf{x}_{i}).$ (4)
Combining Eq. (2), (3), (4), we can get
$p(\mathbf{y}_{i}|\mathbf{e}_{i},\mathbf{x}_{i})\geq p(\mathbf{y}_{i},\mathbf{e}_{i}|\mathbf{x}_{i}).$ (5)
Intuitively, Eq. (5) implies that the learning objective in Eq. (3) can be more deterministic and effective than that in Eq. (2). By considering Eq. (4) and Eq. (5) jointly, we can observe that as the inputs of different problems become more indistinguishable, $p(\mathbf{e}_{i}|\mathbf{x}_{i})$ approaches 0, and the gap between the two learning objectives increases.
Specifically, due to the uncertain correspondence between $\mathbf{x}$ and $\mathbf{e}$, attempting to infer the problem embedding $\mathbf{e}_{i}$ directly from $\mathbf{x}_{i}$ in Eq. (2) is theoretically flawed and likely to introduce noise. Moreover, assuming the output space of $\mathbf{y}$ is of size $Y$, and that of $\mathbf{e}$ is $E$ (even though both are infinite in practice), the output space of Eq. (2) becomes $Y\times E$, which is significantly larger than the output space $Y$ of Eq. (3). This leads to a sparsity of training samples per output configuration in Eq. (2), making the learning less efficient and more prone to underfitting.
@@
IV-B Rank Pool for Problem Feature Embedding
@@
    Structural Sparsity. Mathematical formulations of optimization problems are inherently structured, featuring segmented semantic components, including objective functions, constraints, and optimization variables. However, the corresponding tokens are often distributed sparsely and non-contiguously across the sequence, making it more difficult to retrieve meaningful structure.
@@
    Symbol Sensitivity. Small changes in mathematical symbols can lead to entirely different types of problems. Moreover, global operators such as $\mathrm{min}$, $\sum$, or nested formulations require the structural integrity of the expression to be preserved during token selection.
@@
    Long-Tail Distribution. Standard low-level tokens (e.g., basic operators or variable names) appear frequently but carry limited semantic value. In contrast, high-information global symbols appear less frequently but are crucial for understanding the problem structure. A robust token selection mechanism must retain and emphasize these rare but informative tokens.
Conventional pooling strategies struggle to handle these properties. For instance, MaxPool [55], which selects the most active tokens at each dimension, is highly susceptible to noise and often overemphasizes frequent but low-informative symbols (e.g., repeating input variable names), while potentially breaking mathematical structure by discarding crucial operator relationships. Moreover, MaxPool can fragment tokens, resulting in pooled tokens that are combinations of incomplete parts from multiple original tokens. AvgPool [55], on the other hand, dilutes the semantic strength of key tokens and is insensitive to positional structure, failing to preserve the underlying formulation hierarchy. Although these methods have been widely adopted in other tasks, they demonstrate poor performance when applied to token selection in problem embeddings due to the unique structural and symbolic demands of mathematical optimization representations.
We propose a token selection method based on intra-sequence token similarity scores derived from the problem embedding, referred to as the rank pool. Specifically, given a raw problem embedding sequence of length $S$, we define a similarity score matrix $\mathbf{M}_{\mathrm{SI}}$, where each entry quantifies the similarity between a pair of token vectors. We compute pairwise similarities using PyTorch’s built-in cosine_similarity() function [55], where values approaching 1 indicate high similarity and those near 0 indicate low similarity. Formally, the similarity matrix is defined as
@@
By selecting the most diverse tokens, the rank pool effectively captures unique and informative symbolic elements, particularly for those who are sensitive to small changes in mathematical expressions. Rank pool preserves the integrity of token vectors without fragmenting or averaging their elements, allowing each complete token to potentially convey more of the original information. Additionally, keeping the selected tokens in their original order helps preserve important structural relationships within the expression. In practice, rank pool can be performed only once when the problem is first encountered, and the resulting pooled sequence can be cached for reuse, thereby avoiding repeated calls to the LLM for encoding and pooling. The experimental details are shown in Section V-B.
Input: NN training problems and their corresponding datasets $\boldsymbol{\mathcal{D}}=\{\mathcal{D}_{1},\ldots,\mathcal{D}_{i},\ldots,\mathcal{D}_{N}\}$, the encoder model $\boldsymbol{\theta}_{\mathrm{E}}$, the decoder model $\boldsymbol{\theta}_{\mathrm{D}}$;
@@
5        $\mathbf{x},\mathbf{y}$ $\leftarrow$ batch data from $\mathcal{D}_{i}$;
6        $\mathbf{v}_{\mathrm{joint}}$ $\leftarrow$ cat($\mathbf{x},\mathbf{y}$);
7        $\tilde{\mathbf{v}}_{\mathrm{joint}}$ $\leftarrow$ $f_{\boldsymbol{\theta}_{\mathrm{D}}}(f_{\boldsymbol{\theta}_{\mathrm{E}}}(\mathbf{x}))$;
8        $\mathbf{v}_{\mathrm{val}},\mathbf{v}_{\mathrm{type}}$ $\leftarrow$ split $\mathbf{v}_{\mathrm{joint}}$;
9        $\tilde{\mathbf{v}}_{\mathrm{val}},\tilde{\mathbf{v}}_{\mathrm{type}}$ $\leftarrow$ split $\tilde{\mathbf{v}}_{\mathrm{joint}}$;
10        $\mathcal{L}_{\mathrm{real}}$ $\leftarrow$ $\|\mathbf{v}_{\mathrm{val}}-\tilde{\mathbf{v}}_{\mathrm{val}}\|^{2}$;
11        $\mathcal{L}_{\mathrm{class}}$ $\leftarrow$ $\mathrm{CrossEntropyLoss}(\mathbf{v}_{\mathrm{type}},\tilde{\mathbf{v}}_{\mathrm{type}})$;
12        $\mathcal{L}$ $\leftarrow$ $\mathcal{L}_{\mathrm{real}}+\mathcal{L}_{\mathrm{class}}$;